If log3(x+1)-log3(x-1)=2, then x=A/8 where A=
Answer by stanbon(48535)

If log3(x+1)-log3(x-1)=2, then x=A/8 where A=
-----------------
The base is "3, so:
log[(x+1)/(x-1)] = 2
[(x+1)/(x-1)] = 3^2
x+1 = 9x-9
8x = 10
x = 5/4
----------
If x = A/8
then (5/4) = A/8
A = 10
=============
Cheers,
Stan H.